how many ways are there to form a group of 5 students out of a class of 25?
Answer by jim_thompson5910(35256)


C(n,r) = (n!)/(r!(n-r)!)

C(25,5) = (25!)/(5!*(25-5)!)

C(25,5) = (25!)/(5!*20!)

C(25,5) = (25*24*23*22*21*20!)/(5!*20!)

C(25,5) = (25*24*23*22*21)/(5!)

C(25,5) = (25*24*23*22*21)/(5*4*3*2*1)

C(25,5) = (6375600)/(120)

C(25,5) = 53130

So there are 53130 ways.